# A mathematical theory of topological invariants of quantum spin systems

### v2

@@ 2.1 Locality and (pre-)cosheaves
where $\alpha=\iota_{U\wedge V,U}-\iota_{U\wedge V,V}$ and $\beta=\iota_{U,U\vee V}+\iota_{V,U\vee V}$.
The following proof is essentially a restatement of the proof of Proposition 1.3 in [16], adapted to the site $L$:
Proof.
We must show that for any $U\in X$ and every covering $U_{1}\vee\ldots\vee U_{n}=U$, the sequence of vector spaces
@@
is exact, where $\alpha=\sum_{i<j}\iota_{U_{i}\wedge U_{j},U_{i}}-\iota_{U_{i}\wedge U_{j},U_{j}}$ and $\beta=\sum_{i}\iota_{U_{i},U}$. Exactness of the last three terms follows from an easy induction and the associativity of the join operation. For exactness of the first three terms of the sequence, we also proceed by induction: suppose the result holds for all covers of cardinality $n-1$, and let $U_{1},\ldots,U_{n}\in L$ be a cover of $U$. Suppose $(s_{1},\ldots,s_{n})\in\bigoplus_{i=1}^{n}{\mathfrak{F}}(U_{i})$ lies in the kernel of $\beta$. Let $V:=\bigvee_{i=1}^{n-1}U_{i}$. An application of (3) with $U_{n}$ and $V$ shows that $s_{n}=\iota_{U_{n}\cap V,U}(t)$ for some $t\in V\wedge U_{n}=(U_{1}\wedge U_{n})\vee...\vee(U_{n-1}\vee U_{n})$, and by right-exactness of (4) this shows that $s_{n}=\sum_{i=1}^{n-1}\iota_{U_{i}\cap U_{n},U_{n}}(v_{i})$ for some $v_{i}\in{\mathfrak{F}}(U_{i})$. Finally we write
$(s_{1},\ldots,s_{n})=(s_{1}+w_{1},\ldots,s_{n-1}+w_{n-1},0)+(-w_{1},\ldots,-w_{n-1},s_{n})$
where $w_{i}:=\iota_{U_{i}\cap U_{n},U_{n}}(v_{i})$. Both terms above the first term lies in the image of $\alpha$ by the inductive hypothesis, while the second term equals $\sum_{i=1}^{n-1}\iota_{U_{i}\wedge U_{n},U_{n}}(v_{i})-\iota_{U_{i}\wedge U_{n},U_{i}}(v_{i})$, which evidently is also in the image of $\alpha$. ∎
2.2 DGLA attached to a cover
Let $X$ be a distributive pre-ordered set. Let $W\in X$ and ${\mathfrak{U}}=\{U_{i}\}_{i\in I}$ be a cover of $W$. Let ${\mathfrak{F}}$ be a pre-cosheaf of vector spaces over $X$. The Čech chain complex $C_{\bullet}({\mathfrak{U}},W;{\mathfrak{F}})$ is defined by
$C_{n}({\mathfrak{U}},W;{\mathfrak{F}})=\oplus_{i_{0}<\ldots<i_{n}}{\mathfrak{F}}(U_{i_{0}}\wedge\ldots\wedge U_{i_{n}}),\quad n\geq 0.$
(where some linear order on $I$ has been chosen) with the differential is the Čech differential $\partial:C_{n+1}\rightarrow C_{n}$ given by $\partial=\sum_{j=0}^{n}(-1)^{j}\lambda_{j}$, where $\lambda_{j}$ is the canonical map
${\mathfrak{F}}(U_{i_{0}}\wedge\ldots\wedge U_{i_{n}})\to{\mathfrak{F}}(U_{i_{0}}\wedge\ldots\widehat{U_{i_{j}}}\ldots\wedge U_{i_{n}}).$
Lemma 2.2.
If ${\mathfrak{F}}$ is a coflasque cosheaf of vector spaces over $X$, the homology of the Čech complex is 00 for $n>0$ and ${\mathfrak{F}}(W)$ for $n=0$.
Proof.
Let $C^{aug}({\mathfrak{U}},W;{\mathfrak{F}})=\{C_{\bullet}({\mathfrak{U}},W;{\mathfrak{F}})\rightarrow{\mathfrak{F}}(W)\}$ be the augmented Čech complex.
Proposition 2.1.
Let ${\mathfrak{F}}$ be a local Lie algebra over $X$. Then for any $W\in X$ and any cover ${\mathfrak{U}}$ of $W$ the 1-shifted augmented Čech complex $C^{aug}_{\bullet+1}({\mathfrak{U}},W;{\mathfrak{F}})$ has a natural structure of a non-negatively graded acyclic DGLA.
Proof.
Let ${\mathfrak{U}}$ be a cover of $W$ indexed by $I$. Let $V$ be a vector space with a basis $e_{i}$, $i\in I$ and let $f^{i},i\in I$, be the dual basis of $V^{*}$. The 1-shifted augmented Čech complex of ${\mathfrak{F}}$ with respect to ${\mathfrak{U}}$ is naturally identified with a sub-complex of the DGLA $({\mathfrak{F}}(W)\otimes\Lambda^{\bullet}V,\partial)$, where $\partial$ is contraction with $\sum_{i}f^{i}$. It is easy to check that this sub-complex is closed with respect to the graded Lie bracket thanks to Property I. By Lemma 2.2, the resulting DGLA is acyclic. ∎
Proposition 2.2.
For a fixed $W\in X$, the map which sends a local Lie algebra ${\mathfrak{F}}$ and a cover ${\mathfrak{U}}$ of $W$ to the DGLA $C^{aug}_{\bullet+1}({\mathfrak{U}},W;{\mathfrak{F}})$ is functorial in both ${\mathfrak{F}}$ and ${\mathfrak{U}}$.
Proof.
Functoriality in ${\mathfrak{F}}$ is clear. Functoriality in ${\mathfrak{U}}$ is written Čech component-wise for ${\mathsf{p}}^{\mathfrak{V}}\in C_{k}({\mathfrak{V}},W;{\mathfrak{F}})$:
$(\phi_{*}{\mathsf{p}}^{\mathfrak{V}}_{k})_{i_{0},\dots,i_{k}}:=\sum_{j_{0}\in\phi^{-1}(i_{0})}\cdots\sum_{j_{k}\in\phi^{-1}(i_{k})}({\mathsf{p}}^{\mathfrak{V}}_{k})_{j_{0},\dots,j_{k}},$
for $i_{0}<i_{2}<\cdots<i_{k}$. ∎
We will call $C^{aug}_{\bullet+1}(-,W;-)$ the Čech functor. We will need some variants of the Čech functor. First, we can define a graded local Lie algebra over a distributive pre-ordered set $(X,\leq)$ in an obvious manner. The construction of the acyclic DGLA $C^{aug}_{\bullet+1}({\mathfrak{U}},W;{\mathfrak{F}})$ works in this case as well, except that it may have components in negative degrees.
Second, we define a pointed DGLA as a DGLA equipped with a distinguished central cycle of degree $-2$ (which we call the curvature). A morphism of pointed DGLAs is a DGLA morphism which preserves the distinguished central cycle. [^5] The category of pointed DGLAs as defined here is a full subcategory of the category of curved DGLAs as defined in [17]. There, for a curved DGLA with curvature ${\mathsf{B}}$, ${\mathsf{B}}$ is not required to be central and the derivation $\partial$ satisfies $\partial^{2}={\rm ad}_{\mathsf{B}}$. We say that a graded local Lie algebra ${\mathfrak{F}}$ over $(X,\leq)$ with a terminal object $T\in X$ is pointed if it is equipped with a distinguished central element ${\mathsf{B}}\in{\mathfrak{F}}(T)$ of degree $-2$. Morphisms in the category of pointed graded local Lie algebras are required to preserve the distinguished element. Then the DGLA $C^{aug}_{\bullet+1}({\mathfrak{U}},T;{\mathfrak{F}})$ is an acyclic pointed DGLA, the distinguished central cycle being ${\mathsf{B}}\in C^{aug}_{-1}({\mathfrak{U}},T;{\mathfrak{F}})$. Of course, since the DGLA is acyclic, this central cycle is exact. The construction of a pointed DGLA from a pointed graded local Lie algebra and a cover of $T$ is functorial in both arguments.
3 Quantum lattice systems
3.1 Observables and derivations
We use the $\ell^{\infty}$ metric on ${\mathbb{R}}^{n}$, i. e. $d(x,y):=\max_{i=1,\ldots,n}|x_{i}-y_{i}|$. For any $U,V\subset{\mathbb{R}}^{n}$ we write $\operatorname{diam}(U):=\sup_{x,y\in U}d(x,y)$ and $d(U,V):=\inf_{x\in U,y\in V}d(x,y)$. They take values in extended non-negative reals $[0,\infty]$. Thus $\operatorname{diam}(\varnothing)=0$ and $d(U,\varnothing)=\infty$ for any $U\subset{\mathbb{R}}^{n}$. For a nonempty set $U$ and $r\geq 0$ we define $U^{r}:=\{x\in{\mathbb{R}}^{n}:d(x,U)\leq r\}$ while we set $\varnothing^{r}=\varnothing$.
A quantum lattice system consists of a countable subset $\Lambda\subset{\mathbb{R}}^{n}$ (“the lattice”) and a finite-dimensional complex Hilbert space $V_{j}$ for every $j\in\Lambda$. We make the following assumption on the lattice system [^6] In [11] $\Lambda$ was assumed to be a Delone set, i.e. it was required to be uniformly filling and uniformly discrete. These assumptions were imposed on physical grounds. All the results proved in [11] hold under weaker assumptions adopted in this paper. In [8] $\Lambda$ was taken to be ${\mathbb{Z}}^{n}$ for simplicity. : there is a $C_{\Lambda}>0$ such that the number of points of $\Lambda$ in any hypercube of diameter $d$ is bounded by $C_{\Lambda}(d+1)^{n}$.
For any bounded nonempty $X\subset{\mathbb{R}}^{n}$ let ${\mathscr{A}}(X):=\bigotimes_{j\in X\cap\Lambda}\mathrm{Hom}_{{\mathbb{C}}}(V_{j},V_{j})$. For any $X\subset Y$ there is an inclusion ${\mathscr{A}}(X)\hookrightarrow{\mathscr{A}}(Y)$ and the algebras ${\mathscr{A}}(X)$ form a direct system with respect to these inclusions. We extend this direct system to include the empty set by setting ${\mathscr{A}}(\varnothing)={\mathbb{C}}$ and letting the inclusion ${\mathscr{A}}(\varnothing)\hookrightarrow{\mathscr{A}}(X)$ take $\alpha\mapsto\alpha\boldsymbol{1}$. Each ${\mathscr{A}}(X)$ is a finite-dimensional $C^{*}$-algebra with the operator norm, and the inclusions ${\mathscr{A}}(X)\hookrightarrow{\mathscr{A}}(Y)$ preserve this norm. The normed *-algebra of local observables is
[…]
The algebra of quasi-local observables ${\mathscr{A}}$ is the norm-completion of ${\mathscr{A}}_{\ell}$; it is a $C^{*}$-algebra.
@@ 3.1 Observables and derivations
For any bounded $X\subset{\mathbb{R}}^{n}$, define the normalized trace ${\overline{\rm tr}}:{\mathscr{A}}(X)\to{\mathbb{C}}$ as ${\overline{\rm tr}}({\mathcal{A}})=\operatorname{tr}({\mathcal{A}})/{\sqrt{\dim({\mathscr{A}}(X))}}$. For any bounded $X\subset Y$ the partial trace ${\overline{\rm tr}}_{X^{c}}:{\mathscr{A}}(Y)\to{\mathscr{A}}(X)$ is uniquely specified by the condition ${\overline{\rm tr}}_{X^{c}}({\mathcal{A}}\otimes{\mathcal{B}})={\overline{\rm tr}}({\mathcal{A}}){\mathcal{B}}$ for any ${\mathcal{A}}\in{\mathscr{A}}(Y\backslash X)$ and ${\mathcal{B}}\in{\mathscr{A}}(X)$. Besides forming a direct system with respect to inclusions, the spaces ${\mathscr{A}}(X)$ are also an inverse system with respect to the partial trace. ${\overline{\rm tr}}$ extends to a normalized positive linear functional on ${\mathscr{A}}$, i.e. a state. We say that ${\mathcal{A}}\in{\mathscr{A}}$ is traceless if ${\overline{\rm tr}}({\mathcal{A}})=0$. The space of traceless anti-hermitian elements of ${\mathscr{A}}(X)$ will be denoted ${{\mathfrak{d}}_{l}}(X)$. ${{\mathfrak{d}}_{l}}(X)$ is a real Lie algebra with respect to the commutator. The Lie algebras ${{\mathfrak{d}}_{l}}(X)$ form a direct system over the directed set of bounded subsets of ${\mathbb{R}}^{n}$, and its limit will be denoted ${{\mathfrak{d}}_{l}}$. Equivalently, ${{\mathfrak{d}}_{l}}$ is the Lie algebra of traceless anti-hermitian elements of ${\mathscr{A}}_{\ell}$. Note that ${\mathscr{A}}_{\ell}={\mathbb{C}}{\mathbf{1}}\oplus({{\mathfrak{d}}_{l}}\otimes{\mathbb{C}})$.
Definition 3.1.
@@
where $(k_{1},\ldots,k_{n})$, $(\ell_{1},\ldots,\ell_{n})$, and $(m_{1},\ldots,m_{n})$ are $n$-tuples of integers. We write ${\mathbb{B}}_{n}$ for the set of all bricks in ${\mathbb{R}}^{n}$.
The set of bricks exhausts the collection of bounded subsets of ${\mathbb{R}}^{n}$ in the sense that any bounded subset is contained in a brick. In addition, the set of bricks satisfies the following regularity property:
Lemma 3.1.
For any $j\in{\mathbb{R}}^{n}$ we have
$\sum_{Y\in{\mathbb{B}}_{n}}(1+\operatorname{diam}(Y)+d(Y,j))^{-2n-2}$ $\leq\frac{\pi^{4}4^{n}(n+1)^{2}}{36}$
@@
Any pair of points $x,y\in{\mathbb{Z}}^{n}$ specifies a brick with $x$ and $y$ on opposing corners, and any brick can be specified this way (not uniquely). With $X$ the brick corresponding to $x$ and $y$ it is easy to see that $\max(d(x,j),d(y,j))\leq\operatorname{diam}(X)+d(X,j)$, and so $(1+d(x,j))(1+d(y,j))\leq(1+\operatorname{diam}(X)+d(X,j))^{2}$. Thus we have
$\sum_{Y\in{\mathbb{B}}_{n}}(1+\operatorname{diam}(Y)+d(Y,j))^{-2n-2}$ $\leq\sum_{\begin{subarray}{c}x,y\in{\mathbb{Z}}^{n}\end{subarray}}(1+d(x,j))^{-n-1}(1+d(y,j))^{-n-1}$
$=\left(\sum_{x\in{\mathbb{Z}}^{n}}(1+d(x,j))^{-n-1}\right)^{2},$
and it remains only to bound the above sum. Let $f(k):=(1+k)^{-n-1}$ and $g(k):=\#({\mathbb{Z}}^{n}\cap B_{k}(j))\leq(1+2k)^{n}$. Using summation by parts we have
$\sum_{j\in\Lambda}(1+d(x,j))^{-n-1}$ $\leq\sum_{k\geq 0}f(k)(g(k+1)-g(k))$
$=\lim_{k\to\infty}f(k)g(k)-\sum_{k\geq 0}g(k)(f(k+1)-f(k)).$
It is easy to check that $f(k)g(k)\to 0$ and that $-(f(k+1)-f(k))\leq(n+1)(1+k)^{-n-2}$, and so
$\sum_{j\in\Lambda}(1+d(x,j))^{-n-1}$ $\leq 2^{n}(n+1)\sum_{k\geq 0}(1+k)^{-2}$
$\leq\frac{\pi^{2}2^{n}(n+1)}{6}$

### v3

@@ 2.1 Observables and derivations
The algebra of quasi-local observables $\mathscr{A}$} is the norm-completion of $\mathscr{A}$$_{\ell}$; it is a ${\mathbb{Z}}_{2}$-graded $C^{*}$-algebra.
Example 3.
For systems of fermionic particles on a lattice, ${\mathscr{A}}_{x}$ as well as ${\mathscr{A}}(X)$ for any bounded $X\subseteq{\mathbb{R}}^{n}$ are complex Clifford algebras. For simplicity, assume that all $H_{x}$, $x\in\Lambda$, are isomorphic to the same Hilbert space $H$ of dimension $N$. Choose an orthonormal basis in $H$. Then ${\mathscr{A}}(X)$ is generated as a ${\mathbb{Z}}_{2}$-graded *-algebra by odd local observables $c_{\alpha,x}$, $\alpha\in\{1,\ldots,N\}$, $x\in X\cap\Lambda$. The relations are
$c_{\alpha,x}c_{\beta,y}+c_{\beta,y}c_{\alpha,x}$ $=0,$ $\forall\alpha,\beta,\forall x,y\in X\cap\Lambda,$ (3)
$c_{\alpha,x}c^{*}_{\beta,y}+c^{*}_{\beta,y}c_{\alpha,x}$ $=\delta_{\alpha\beta}\delta_{xy},$ $\forall\alpha,\beta,\forall x,y\in X\cap\Lambda.$ (4)
The algebra of quasi-local observables $\mathscr{A}$} for systems of this type is known as a CAR algebra (the algebra of canonical anti-commutation relations) [11, 1].
For any bounded $X\subseteq{\mathbb{R}}^{n}$, define the normalized even trace ${\overline{\rm tr}}:{\mathscr{A}}(X)\to{\mathbb{C}}$ as ${\overline{\rm tr}}({\mathcal{A}})=\operatorname{tr}({\mathcal{A}})/{\sqrt{\dim({\mathscr{A}}(X))}}$. For any bounded $X\subseteq Y$ the partial trace ${\overline{\rm tr}}_{X^{c}}:{\mathscr{A}}(Y)\to{\mathscr{A}}(X)$ is uniquely specified by the condition
${\overline{\rm tr}}_{X^{c}}({\mathcal{A}}\otimes{\mathcal{B}})={\overline{\rm tr}}({\mathcal{A}}){\mathcal{B}}$ (5)
@@
We say that ${\mathcal{A}}\in{\mathscr{A}}$ is traceless if ${\overline{\rm tr}}({\mathcal{A}})=0$. The space of traceless anti-hermitian even elements of ${\mathscr{A}}(X)$ will be denoted by ${{\mathfrak{d}}_{l}}(X)$. ${{\mathfrak{d}}_{l}}(X)$ is a real Lie algebra with respect to the commutator. Note that ${{\mathfrak{d}}_{l}}(X)=0$ if $X\cap\Lambda=\emptyset$. The Lie algebras ${{\mathfrak{d}}_{l}}(X)$ form a direct system over the directed set of bounded subsets of ${\mathbb{R}}^{n}$, and its limit will be denoted ${\mathfrak{d}}_{l}$}. Equivalently, ${\mathfrak{d}}_{l}$} is the Lie algebra of traceless anti-hermitian even elements of $\mathscr{A}$$_{\ell}$. Note that ${\mathscr{A}}_{\ell}^{even}={\mathbb{C}}{\mathbf{1}}\oplus({{\mathfrak{d}}_{l}}\otimes{\mathbb{C}})$.
So far, our discussion has been completely standard and mostly follows the monograph [10, 11]. A point of departure is the way we treat the derivations of $\mathscr{A}$}. In the $C^{*}$-algebraic approach, derivations represent generators of continuous symmetries of a quantum lattice system. For example, a Hamiltonian is a generator of the time-translation symmetry and is described by a derivation. For quantum lattice systems, derivations of physical interest have the following form [11]:
Definition 4.
Let $\mathcal{P}(\Lambda)$ denote the set of finite subsets of $\Lambda$. An interaction is a function $\Phi:\mathcal{P}(\Lambda)\to{{\mathfrak{d}}_{l}}$ such that $\Phi(X)\in{{\mathfrak{d}}_{l}}(X)$ for each $X\in\mathcal{P}(\Lambda)$ and such that for every $x\in\Lambda$ we have
$\sum_{\begin{subarray}{c}X\in\mathcal{P}(\Lambda)\\ X\ni x\end{subarray}}\|\Phi(X)\|<\infty.$ (6)
An interaction determines a derivation ${\mathsf{F}}_{\Phi}$ defined by
${\mathsf{F}}_{\Phi}:{\mathcal{A}}\mapsto\sum_{X\in{\mathcal{P}}(\Lambda)}[\Phi(X),{\mathcal{A}}],\quad{\mathcal{A}}\in{\mathscr{A}}.$ (7)
This is a (possibly unbounded) operator on $\mathscr{A}$} whose domain is taken to be $\mathscr{A}$$_{\ell}$.
Physically interesting interactions are typically supported on an infinite subset of ${\mathcal{P}}(\Lambda)$, as we see in the examples below.
Example 5.
An on-site action of $U(1)$ on a lattice system $(\Lambda,\{V_{x}\}_{x\in\Lambda})$ is a collection of continuous (and therefore smooth) homomorphisms $\rho_{x}:U(1)\rightarrow U(V_{x})^{even}$. An on-site action of $U(1)$ on $(\Lambda,\{V_{x}\}_{x\in\Lambda})$ gives rise to a homomorphism from $U(1)$ to the automorphism group of $\mathscr{A}$} via $g\mapsto\otimes_{x\in\Lambda}{\rm Ad}_{\rho_{x}(g)}$. The corresponding generator $\mathsf{Q}$} is a derivation of the form
${\mathsf{Q}}:{\mathcal{A}}\mapsto\sum_{x\in\Lambda}[{\mathsf{q}}_{x},{\mathcal{A}}],\quad{\mathcal{A}}\in{\mathscr{A}},$ (8)
where ${\mathsf{q}}_{x}\in{{\mathfrak{d}}_{l}}(\{x\})$ is the traceless part of the generator of $\rho_{x}$. The domain of $\mathsf{Q}$} is dense because it contains $\mathscr{A}$$_{\ell}$. It is easy to see that for an infinite $\Lambda$, $\mathsf{Q}$} is unbounded even if we require the norms of ${\mathsf{q}}_{x}$ to be bounded uniformly in $x$.
Example 6.
Systems of fermionic particles (Example 3) have a natural on-site $U(1)$ action (particle number symmetry). It is defined by saying that the weight-$\ell$ component of $V_{x}$ is the degree-$\ell$ component of the exterior algebra $\wedge H_{x}$. The generator of this $U(1)$ action has the form (8) with ${\mathsf{q}}_{x}$ being the grading operator on site $x$. More explicitly, suppose $H_{x}\simeq{\mathbb{C}}^{N}$ for all $x\in\Lambda$. Choose an orthonormal basis for $H$, then $\mathscr{A}$$_{\ell}$ is generated as a ${\mathbb{Z}}_{2}$-graded *-algebra by odd local observables $c_{\alpha,x}\in{\mathscr{A}}_{x}$, $\alpha\in\{1,\ldots,N\}$, $x\in\Lambda$, with the relations as in (3). The charge derivation has the form (8) with ${\mathsf{q}}_{x}=\sqrt{-1}\sum_{\alpha}\left(c^{*}_{\alpha,x}c_{\alpha,x}-\frac{1}{2}\right)$. The term 1/2-1/2 in parentheses is needed to make ${\mathsf{q}}_{x}$ traceless, while the overall $\sqrt{-1}$ makes it anti-hermitian.
The brick decomposition
The specification (7) of a derivation in terms of an interaction $\Phi$ is many-to-one. For instance, for any $X,Y\supset Z\in\mathcal{P}(\Lambda)$ and ${\mathsf{a}}\in{{\mathfrak{d}}_{l}}(Z)$ we may redefine the interaction by setting $\Phi^{\prime}(X)=\Phi(X)+{\mathsf{a}}$ and $\Phi^{\prime}(Y)=\Phi(Y)-{\mathsf{a}}$, yielding the same derivation. It is useful to have a way of describing unbounded derivations that does not suffer from this ambiguity. This is the purpose of the brick decomposition [19] which we review now.
Definition 7.
@@
where $(\ell_{1},\ldots,\ell_{n})$ and $(m_{1},\ldots,m_{n})$ are $n$-tuples of integers. We write ${\mathbb{B}}_{n}$ for the set of all bricks in ${\mathbb{R}}^{n}$.
The set of bricks ${\mathbb{B}}_{n}$ is a poset with respect to inclusion. In fact, it is a locally-finite distributive lattice. We denote the meet and join in ${\mathbb{B}}_{n}$ by $\wedge$ and $\vee$, respectively. The set of bricks exhausts the collection of bounded subsets of ${\mathbb{R}}^{n}$ in the sense that any bounded subset is contained in a brick. ${\mathbb{B}}_{n}$ satisfies the following regularity property:
Lemma 8.
For any $z\in{\mathbb{R}}^{n}$ we have
$\sum_{Y\in{\mathbb{B}}_{n}}(1+\operatorname{diam}(Y)+d(Y,z))^{-2n-2}\leq\frac{\pi^{4}4^{n}(n+1)^{2}}{36}$ (10)
@@
Any pair of points $x,y\in{\mathbb{Z}}^{n}$ specifies a brick with $x$ and $y$ on opposing corners, and any brick can be specified this way (not uniquely). With $X$ the brick corresponding to $x$ and $y$ it is easy to see that $\max(d(x,z),d(y,z))\leq\operatorname{diam}(X)+d(X,z)$, and so $(1+d(x,z))(1+d(y,z))\leq(1+\operatorname{diam}(X)+d(X,z))^{2}$. Thus we have
$\sum_{Y\in{\mathbb{B}}_{n}}(1+\operatorname{diam}(Y)+d(Y,z))^{-2n-2}$ $\leq\sum_{\begin{subarray}{c}x,y\in{\mathbb{Z}}^{n}\end{subarray}}(1+d(x,z))^{-n-1}(1+d(y,z))^{-n-1}$
$=\left(\sum_{x\in{\mathbb{Z}}^{n}}(1+d(x,z))^{-n-1}\right)^{2},$ (11)
and it remains only to bound the above sum. Let $f(k):=(1+k)^{-n-1}$ and $g(k):=\#({\mathbb{Z}}^{n}\cap B_{k}(z))\leq(1+2k)^{n}$. Using summation by parts we have
$\sum_{x\in{\mathbb{Z}}^{n}}(1+d(x,z))^{-n-1}$ $\leq\sum_{k\geq 0}f(k)(g(k+1)-g(k))$
$=\lim_{k\to\infty}f(k)g(k)-\sum_{k\geq 0}g(k)(f(k+1)-f(k)).$ (12)
It is easy to check that $f(k)g(k)\to 0$ and that $(f(k+1)-f(k))\leq(n+1)(1+k)^{-n-2}$, and so
$\sum_{x\in{\mathbb{Z}}^{n}}(1+d(x,z))^{-n-1}$ $\leq 2^{n}(n+1)\sum_{k\geq 0}(1+k)^{-2}$
$\leq\frac{\pi^{2}2^{n}(n+1)}{6}$ (13)